OpenStudy (anonymous):

I Neeed help pleasee What is the sine value of 5 pi over 3? negative square root 3 over 2 square root 3 over 2 negative 1 over 2 1 over 2

OpenStudy (anonymous):

\[\sin \frac{ 5 \pi }{ 3 }= \sin \left(2 \pi-\frac{ \pi }{ 3 } \right)=-\sin \frac{ \pi }{ 3 }=?\]

OpenStudy (anonymous):

find \(\frac{5\pi}{3}\) on the unit circle on the last page of the attached cheat sheet the first coordinate is cosine, the second is sine

OpenStudy (anonymous):

I found it but I see two answers,how Will I go about it to find the right answer

OpenStudy (anonymous):

there are not two answers, only one if you mean the point has a first and a second coordinate, sine is the second coordinate cosine is the first

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

yw